OpenStudy (jacksonjrb):

Solve for x in terms of y. y=x^2-4x+3

OpenStudy (solomonzelman):

ok, I will show a similar problem.

OpenStudy (jacksonjrb):

ok

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle y=2x^2-3x+5 }\) \(\large\color{black}{ \displaystyle 0=2x^2-3x+5-y }\) so a=2 b=-3 c=(5-y) \(\large\color{black}{ \displaystyle x=\frac{-(-3)\pm\sqrt{(-3)^2-4(2)(5-y)}}{2(2)} }\) \(\large\color{black}{ \displaystyle x=\frac{3\pm\sqrt{9-8(5-y)}}{4} }\) \(\large\color{black}{ \displaystyle x=\frac{3\pm\sqrt{9-40+8y}}{4} }\) \(\large\color{black}{ \displaystyle x=\frac{3\pm\sqrt{-31+8y}}{4} }\)

OpenStudy (jacksonjrb):

So, i just use quadratic formula?

OpenStudy (solomonzelman):

Yes

OpenStudy (jacksonjrb):

Okay, I'll try it. Thanks

OpenStudy (solomonzelman):

ok

OpenStudy (jacksonjrb):

\[x=\frac{ 4+/- \sqrt{4+4y} }{ 2 }\]

OpenStudy (solomonzelman):

let me check ...

OpenStudy (jacksonjrb):

\[x=2+/-\sqrt{1+y}\]

OpenStudy (jacksonjrb):

Simplifies to that I believe

OpenStudy (solomonzelman):

y=x^2-4x+3 0=x^2-4x+3-y \(\large\color{black}{ \displaystyle x=\frac{4\pm\sqrt{(-4)^2-4(1)(3-y)}}{2} }\) \(\large\color{black}{ \displaystyle x=\frac{4\pm\sqrt{16-4(3-y)}}{2} }\) \(\large\color{black}{ \displaystyle x=\frac{4\pm\sqrt{4(4-(3-y))}}{2} }\) \(\large\color{black}{ \displaystyle x=\frac{4\pm2\sqrt{4-(3-y)}}{2} }\) \(\large\color{black}{ \displaystyle x=2\pm\sqrt{4-(3-y)} }\) \(\large\color{black}{ \displaystyle x=2\pm\sqrt{4-3+y} }\) \(\large\color{black}{ \displaystyle x=2\pm\sqrt{1+y} }\) Very Nice!

OpenStudy (solomonzelman):

Good job, keep that up:)

OpenStudy (jacksonjrb):

Thanks for the help

OpenStudy (solomonzelman):

you have done it all yourself, i just pointed:)

OpenStudy (jacksonjrb):

There's another problem that involves fractions though

OpenStudy (solomonzelman):

ok, ...

OpenStudy (jacksonjrb):

Little confused

OpenStudy (jacksonjrb):

\[y=\frac{ x^2+1 }{ x^2-1 }\]

OpenStudy (solomonzelman):

you want x in terms of y?

OpenStudy (jacksonjrb):

yes

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle y=\frac{x^2+1}{x^2-1} }\) \(\large\color{black}{ \displaystyle y=\frac{x^2-1+2}{x^2-1} }\) \(\large\color{black}{ \displaystyle y=\frac{x^2-1}{x^2-1}+\frac{2}{x^2-1} }\) \(\large\color{black}{ \displaystyle y=1+\frac{2}{x^2-1} }\) can you get it from there?

OpenStudy (jacksonjrb):

Why did you add 2?

OpenStudy (solomonzelman):

I didn't, I rewrote 1 as 2-1

OpenStudy (solomonzelman):

-1+2, is same as 1

OpenStudy (jacksonjrb):

Oh, I see now

OpenStudy (jacksonjrb):

I think I can finish it

OpenStudy (solomonzelman):

ok, go ahead, but take your time...

OpenStudy (jacksonjrb):

\[x=\sqrt{\frac{ 2 }{ y-1 }+1}\]

OpenStudy (solomonzelman):

yes (I think though you need the \(\pm\) there)

OpenStudy (jacksonjrb):

You're right

OpenStudy (jacksonjrb):

Thanks

OpenStudy (solomonzelman):

But, I mean without the \(\pm\) you have still achieved the task. You solved for x in terms of y. You might not need ± because (-)² or (+)² is all 1.

OpenStudy (solomonzelman):

yw in any case.